OpenStudy (xo.msml):

HELP PLEASEEE Variable a is 4 more than variable b. Variable a is also 1 less than b. Which of the following pairs of equations best models the relationship between a and b? 1.a = b + 4 a = b − 1 2. a = 4b a = b + 1 3. b = a + 4 b = a − 1 4. b = 4a b = a + 1

OpenStudy (mjdennis):

Try this: start with 'b' by itself: b Then write "4 more than b" in math symbols. Does that start you off well?

OpenStudy (skullpatrol):

$$\Huge \text{a is 4 more than b}$$ $$\Huge a = 4 + b$$

OpenStudy (skullpatrol):

Recall: a + b = b + a let a = 4 therefore 4 + b = b + 4.

OpenStudy (skullpatrol):

$$\Huge\text{ a is 1 less than b}$$ $$\Huge a = b – 1$$

OpenStudy (skullpatrol):

Recall "is" means "equals"